jason is 22 years older than his sister Kate. In 14 years, he will be twice as old as Kate. How old is each of them now?

Answer by ikleyn(52792)

.
J = K + 22.             (1)


In 14 years:


J + 14 = 2*(K + 14)     (2)


Substitute (1) into (2):


(K+22) + 14 = 2*(K+14)


and solve:


K + 36 = 2K + 28

36 - 28 = 2K - K

K = 8     is the Kate's age     ANSWER

J = K + 22 = 8 + 22 = 30   is the Jason's age.      ANSWER

Solved.